### v2

# Linear equations and multiplicative polynomial equations in infinitely many variables

Abstract.
This paper describes infinite sets of linear and polynomial equations in infinitely many variables with the property that the existence of a solution or even an approximate solution for every finite subset of the equations implies the existence of a solution for the infinite set of equations.
Key words and phrases:
Systems of polynomial equations, systems of linear equations, approximate solutions, infinitely many variables, finite intersection property, Dirichlet series.
2000 Mathematics Subject Classification:
12D10, 12E12, 15A06, 40H05, 46A45, 54B10, 54C30
MBN supported in part by PSC-CUNY Research Award Program grant 66197-00 54.
1. Finitely many implies infinite many
In mathematics there are theorems asserting that, for certain classes of equations, if every finite subset of an infinite set of the equations has a solution, then the infinite set of equations has a solution. For example, if every finite subset of an infinite set of linear equations in $n$ variables (that is, equations of the form $\sum_{j=1}^{n}a_{i,j}x_{j}=b_{i}$) has a solution, then the infinite set of linear equations in $n$ variables has a solution. For infinite sets of linear equations in infinitely many variables, in which each equation contains only finitely many variables (that is, equations of the form $\sum_{j=1}^{\infty}a_{i,j}x_{j}=b_{i}$ such that, for each $i$, we have $a_{i,j}\neq 0$ for only finitely many $j$), a “finitely solvable implies infinitely solvable” theorem is also true. For a survey of related results, see Nathanson [4].
There are analogous results for linear equations that contain infinitely many variables. The pair $(p,q)$ is Hölder conjugate or, simply, conjugate if $p>1$ and $q>1$ are real numbers such that
$\frac{1}{p}+\frac{1}{q}=1$
or if $p=1$ and $q=\infty$. For $1\leq p<\infty$, let
$\ell^{p}=\left\{\mathbf{a}=\left(a_{j}\right)_{j=1}^{\infty}:\|\mathbf{a}\|_{p}=\left(\sum_{j=1}^{\infty}|a_{j}|^{p}\right)^{1/p}<\infty\right\}$
and, for $p=\infty$, let
=$\ell^{\infty}=\left\{\mathbf{a}=\left(a_{j}\right)_{j=1}^{\infty}:\|\mathbf{a}\|_{\infty}=\sup\{|a_{j}|:j=1,2,3,\ldots\}<\infty\right\}$
be the classical Lebesque spaces of infinite sequences of real or complex numbers.
Recall Hölder’s inequality: If $(p,q)$ is a conjugate pair and
$\mathbf{a}=\left(a_{j}\right)_{j=1}^{\infty}\in\ell^{p}\operatorname{\qquad\text{and}\qquad}\mathbf{x}=\left(x_{j}\right)_{j=1}^{\infty}\in\ell^{q}$
then
$\mathbf{a}\mathbf{x}=\left(a_{j}x_{j}\right)_{j=1}^{\infty}\in\ell^{1}$
and
$\sum_{j=1}^{\infty}|a_{j}x_{j}|=\left\|\mathbf{a}\mathbf{x}\right\|_{1}\leq\left\|\mathbf{a}\right\|_{p}\left\|\mathbf{x}\right\|_{q}.$
Thus, the infinite series
$(\mathbf{a},\mathbf{x})=\sum_{j=1}^{\infty}a_{j}x_{j}$
converges absolutely and it makes sense to ask if there is a solution $\mathbf{x}\in\ell^{q}$ of the linear equation in infinitely many variables
(1) $L(\mathbf{x})=(\mathbf{a},\mathbf{x})=\sum_{j=1}^{\infty}a_{j}x_{j}=b.$
We retain the usual ambiguity between sequences of numbers and sequences of variables. An exact solution or, simply, a solution of equation (1) is a sequence $\mathbf{x}=(x_{j})_{j=1}^{\infty}\in\ell^{q}$ such that $L(\mathbf{x})=b$. Equation (1) has an approximate solution if, for every $\varepsilon>0$, there exists $\mathbf{x}_{\varepsilon}\in\ell^{q}$ such that
$\left|L(\mathbf{x}_{\varepsilon})-b\right|<\varepsilon.$
Let $\mathbf{N}=\{1,2,3,\ldots\}$ be the set of positive integers and $\mathbf{N}^{k}$ the set of $k$-tuples of positive integers. Extending a classical result of F. Riesz [5, pp. 61–62] (see also Banach [2, p. 47]), Abian and Eslami [1] proved the following beautiful theorem.
Theorem 1.
Let $(p,q)$ be a conjugate pair with $p>1$ and $q>1$ and let $\mathbf{a}_{i}=\left(a_{i,j}\right)_{j=1}^{\infty}\in\ell^{p}$ for all $i\in\mathbf{N}$. Consider the linear equation in infinitely many variables
$L_{i}(\mathbf{x})=\sum_{j=1}^{\infty}a_{i,j}x_{j}=b_{i}.$
Let $M>0$. If, for every finite subset $S$ of $\mathbf{N}$, there is a sequence $\mathbf{x}_{S}\in\ell^{q}$ such that $\|\mathbf{x}_{S}\|_{q}\leq M$ and $L_{i}(\mathbf{x}_{S})=b_{i}$ for all $i\in S$, then there is a sequence $\mathbf{x}\in\ell^{q}$ such that $\|\mathbf{x}\|_{q}\leq M$ and $L_{i}(\mathbf{x})=b_{i}$ for all $i\in\mathbf{N}$.
Abian and Eslami proved their theorem only for $p>1$ and for countably many equations with real coefficients and real solutions, but the result also holds for $(p,q)=(1,\infty)$ and for uncountably many equations with complex coefficients and complex solutions. In Appendix A we prove Riesz’s theorem and derive Theorem 1 from it.
It is important to observe that Theorem 1 is not true without the condition that the norm of $\|\mathbf{x}_{S}\|_{q}$ is uniformly bounded by $M$ for all finite sets $S$. The following example is essentially due to Helly [3].
Let $\mathbf{a}=(a_{j})_{j=1}^{\infty}\in\ell^{p}$ with $a_{j}\neq 0$ for all $j\in\mathbf{N}$. For all $i\in\mathbf{N}$, define the sequence $\mathbf{a}_{i}=(a_{i,j})_{j=1}^{\infty}\in\ell^{p}$ as follows:
$a_{i,j}=\begin{cases}0&\text{if $j<i$}\\ a_{i}&\text{if $j\geq i$.}\end{cases}$
For all $i\in\mathbf{N}$, consider the linear equation in infinitely many variables
$L_{i}(\mathbf{x})=\sum_{j=1}^{\infty}a_{i,j}x_{j}=\sum_{j=i}^{\infty}a_{j}x_{j}=1.$
We obtain the “triangular” system of equations
$\begin{array}[]{rl}a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}+\cdots+a_{i}x_{i}+a_{i+1}x_{i+1}+\cdots&=1\\ a_{2}x_{2}+a_{3}x_{3}+\cdots+a_{i}x_{i}+a_{i+1}x_{i+1}+\cdots&=1\\ a_{3}x_{3}+\cdots+a_{i}x_{i}+a_{i+1}x_{i+1}+\cdots&=1\\ \vdots&\\ a_{i}x_{i}+a_{i+1}x_{i+1}+\cdots&=1\\ a_{i+1}x_{i+1}+\cdots&=1\\ \end{array}$
For all $i\geq 1$, if $L_{i}(\mathbf{x})=L_{i+1}(\mathbf{x})=1$, then $a_{i}x_{i}=0$ and so $x_{i}=0$ (because $a_{i}\neq 0$). It follows that the infinite set of equations has no solution. However, every finite subset of these equations is solvable.
Here is one solution of the triangular system. For all $r\in\mathbf{N}$, the sequence $\mathbf{x}_{r}=(x_{r,j})_{j=1}^{\infty}\in\ell^{q}$ defined by
$x_{r,j}=\begin{cases}1/a_{r}&\text{if $j=r$}\\ 0&\text{if $j\neq r$}\end{cases}$
is a solution of the finite set of equations $\{L_{i}(\mathbf{x})=1:i=1,\ldots,r\}$. We have
$\|\mathbf{x}_{r}\|_{q}=\frac{1}{|a_{r}|}\operatorname{\qquad\text{and}\qquad}\lim_{r\rightarrow\infty}\|\mathbf{x}_{r}\|_{q}=\lim_{r\rightarrow\infty}\frac{1}{|a_{r}|}=\infty.$
Thus, the sequence of solutions $(\mathbf{x}_{r})_{r=1}^{\infty}$ is not uniformly bounded in $\ell^{q}$.
More generally, if $\mathbf{x}_{r}$ is any solution of the first $r$ equations of the triangular system, then
$1=\sum_{j=r}^{\infty}a_{j}x_{j}=\sum_{j=1}^{\infty}a_{r,j}x_{j}\leq\|\mathbf{a}_{r}\mathbf{x}_{r}\|_{1}\leq\|\mathbf{a}_{r}\|_{p}\|\mathbf{x}_{r}\|_{q}$
and so
$\|\mathbf{x}_{r}\|_{q}\geq\frac{1}{\|\mathbf{a}_{r}\|_{p}}.$
Because $\mathbf{a}\in\ell^{p}$, we have $\lim_{r\rightarrow\infty}\|\mathbf{a}_{r}\|_{p}=0$ and so the sequence of solutions $(\mathbf{x}_{r})_{r=1}^{\infty}$ is not uniformly bounded in $\ell^{q}$.
The following result extends Theorem 1 to approximately solvable systems of linear equations.
Theorem 2.
Let $(p,q)$ be a conjugate pair. Let $I$ be an infinite set. For all $i\in I$, let $\mathbf{a}_{i}=\left(a_{i,j}\right)_{j=1}^{\infty}\in\ell^{p}$ and consider the linear equation in infinitely many variables
$L_{i}(\mathbf{x})=\sum_{j=1}^{\infty}a_{i,j}x_{j}=b_{i}.$
Let $M>0$. If, for every finite subset $S$ of I and every $\varepsilon>0$, the finite set of linear inequalities
$\left\{|L_{i}(\mathbf{x})-b_{i}|\leq\varepsilon:i\in S\right\}$
has a solution $\mathbf{x}_{S,\varepsilon}\in\ell^{q}$ with $\|\mathbf{x}_{S,\varepsilon}\|_{q}\leq M$, then the infinite set of linear equations
$\left\{L_{i}(\mathbf{x})=b_{i}:i\in I\right\}$
has an exact solution $\mathbf{x}\in\ell^{q}$ with $\|\mathbf{x}\|_{q}\leq M$.
In this paper we generalize Theorem 2 to certain infinite sets of polynomial equations in infinitely many variables and prove that the existence of norm-bounded or even sequentially bounded approximate solutions to all finite subsets of the set of polynomial equations is sufficient to guarantee the existence of an exact solution to the infinite set (Theorems 4 and 5). The theorem on linear equations is a special case of the polynomial results.
Our theorems apply to polynomials with coefficients and solutions in the field of real numbers and also in the field of complex numbers. However, the “finitely many implies infinitely many” paradigm is not true for all subfields of the complex numbers. For example, the paradigm fails for the field $\mathbf{Q}$ of rational numbers. Let $(b_{i})_{i\in I}$ be an infinite sequence of irrational numbers and consider the infinite set of linear equations $\{x_{i}=b_{i}:i\in I\}$. For every $\varepsilon>0$, every finite subset of the equations and, indeed, the infinite set of equations have approximate solutions in $\mathbf{Q}$, but neither the infinite set nor any nonempty finite subset of the equations has an exact solution in $\mathbf{Q}$.
2. Polynomials in infinitely many variables
Let $D$ be a positive integer, let $d\in\{1,2,\ldots,D\}$, and let
$D<q\leq\infty.$
If $q<\infty$, then
$\frac{1}{q/(q-d)}+\frac{1}{q/d}=1$
and so the real numbers $q/(q-d)>1$ and $q/d>1$ form a conjugate pair. For $q=\infty$, we define
$\frac{q}{(q-d)}=1\operatorname{\qquad\text{and}\qquad}\frac{q}{d}=\infty.$
If
$\mathbf{a}=\left(a_{j}\right)_{j=1}^{\infty}\in\ell^{q/(q-d)}$
and
$\mathbf{x}=\left(x_{j}\right)_{j=1}^{\infty}\in\ell^{q}$
then
$\mathbf{x}^{d}=\left(x_{j}^{d}\right)_{j=1}^{\infty}\in\ell^{q/d}.$
From Hölder’s inequality we obtain
$\mathbf{a}\mathbf{x}^{d}=\left(a_{j}x_{j}^{d}\right)_{j=1}^{\infty}\in\ell^{1}$
and
$\sum_{j=1}^{\infty}\left|a_{j}x_{j}^{d}\right|=\left\|\mathbf{a}\mathbf{x}^{d}\right\|_{1}\leq\left\|\mathbf{a}\right\|_{q/(q-d)}\left\|\mathbf{x}^{d}\right\|_{q/d}<\infty.$
Thus, the infinite series
$\left(\mathbf{a},\mathbf{x}^{d}\right)=\sum_{j=1}^{\infty}a_{j}x_{j}^{d}$
converges absolutely.
For all $k\in\{1,2,\ldots,D\}$, let $\mathcal{D}_{k}$ be the set of all $k$-tuples of positive integers whose sum is at most $D$, that is,
(2) $\mathcal{D}_{k}=\left\{(d_{1},\ldots,d_{k})\in\mathbf{N}^{k}:d_{1}+d_{2}+\cdots+d_{k}\leq D\right\}.$
This is a finite set of cardinality $\binom{D}{k}$. The set $\bigcup_{k=1}^{D}\mathcal{D}_{k}$ has cardinality $\sum_{k=1}^{D}\binom{D}{k}=2^{D}-1$. For all
$\Delta=(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}$
and
$J=(j_{1},j_{2},\ldots,j_{k})\in\mathbf{N}^{k}$
we define the monomial
$x_{J}^{\Delta}=x_{j_{1}}^{d_{1}}x_{j_{2}}^{d_{2}}\cdots x_{j_{k}}^{d_{k}}$
of degree
$|\Delta|=d_{1}+d_{2}+\cdots+d_{k}\leq D.$
For example, let $D=3$ and $k\in\{1,2,3\}$. There are $17=2^{3}-$ 1 sets of monomials. For $k=1$ and $J=(j_{1})\in\mathbf{N}^{1}$, we have
$\Delta\in\mathcal{D}_{1}=\{(1),(2),(3)\}\operatorname{\quad\text{and}\quad}x_{J}^{\Delta}\in\left\{x_{j_{1}},\ x_{j_{1}}^{2},\ x_{j_{1}}^{3}\right\}.$
For $k=2$ and $J=(j_{1},j_{2})\in\mathbf{N}^{2}$, we have
$\Delta\in\mathcal{D}_{2}=\{(1,1),(1,2),(2,1)\}\operatorname{\quad\text{and}\quad}x_{J}^{\Delta}\in\left\{x_{j_{1}}x_{j_{2}},\ x_{j_{1}}x_{j_{2}}^{2},\ x_{j_{1}}^{2}x_{j_{2}}\right\}.$
For $k=3$ and $J=(j_{1},j_{2},j_{3})\in\mathbf{N}^{3}$, we have
$\Delta\in\mathcal{D}_{3}=\{(1,1,1)\}\operatorname{\quad\text{and}\quad}x_{J}^{\Delta}=x_{j_{1}}x_{j_{2}}x_{j_{3}}.$
Monomials do not have a unique representation in the form $x_{J}^{\Delta}$. For example, with $D=3$, let $k=2$ and $J=(1,1)\in\mathbf{N}^{2}$. If $\Delta=(1,2)\in\mathcal{D}_{2}$, then $x_{J}^{\Delta}=x_{1}x_{1}^{2}=x_{1}^{3}$. If $\Delta=(2,1)\in\mathcal{D}_{2}$, then $x_{J}^{\Delta}=x_{1}^{2}x_{1}=x_{1}^{3}$. If $D=3$, $k=1$, $J=(1)\in\mathbf{N}^{1}$, and $\Delta=(3)\in\mathcal{D}_{1}$, then $x_{J}^{\Delta}=x_{1}^{3}$. Note that we do not need unique representation of monomials, but we would have unique representation if we considered only $k$-tuples $(j_{1},j_{2},\ldots,j_{k})\in\mathbf{N}^{k}$ such that $j_{1}<j_{2}<\cdots<j_{k}$.)
We consider only polynomials with zero constant term. In this paper, “polynomial” means polynomial with zero constant term. By moving constants from one side of an equation to the other side, it is always possible to transform an equation that includes polynomials with nonzero constant terms to an equivalent equation in which no polynomial has a nonzero constant term.
A polynomial of degree at most $D$ in infinitely many variables $\mathbf{x}=(x_{j})_{j=1}^{\infty}$ is a formal power series of the form
(3) $P(\mathbf{x})=\sum_{k=1}^{D}\sum_{\Delta\in\mathcal{D}_{k}}\sum_{J\in\mathbf{N}^{k}}a_{\Delta,J}\ x_{J}^{\Delta}$
with real or complex coefficients $a_{\Delta,J}$. For example, a polynomial of degree at most 3 in infinitely many variables is of the form
$P(\mathbf{x})$ $=\sum_{j_{1}=1}^{\infty}a_{(1),(j_{1})}x_{j_{1}}+\sum_{j_{1}=1}^{\infty}a_{(2),(j_{1})}x_{j_{1}}^{2}+\sum_{j_{1}=1}^{\infty}a_{(3),(j_{1})}x_{j_{1}}^{3}$
$+\sum_{j_{1}=1}^{\infty}\sum_{j_{2}=1}^{\infty}a_{(1,1),(j_{1},j_{2})}x_{j_{1}}x_{j_{2}}+\sum_{j_{1}=1}^{\infty}\sum_{j_{2}=1}^{\infty}a_{(1,2),(j_{1},j_{2})}x_{j_{1}}x_{j_{2}}^{2}$
$\quad+\sum_{j_{1}=1}^{\infty}\sum_{j_{2}=1}^{\infty}a_{(2,1),(j_{1},j_{2})}x_{j_{1}}^{2}x_{j_{2}}+\sum_{j_{1}=1}^{\infty}\sum_{j_{2}=1}^{\infty}\sum_{j_{3}=1}^{\infty}a_{(1,1,1),(j_{1},j_{2},j_{3})}x_{j_{1}}x_{j_{2}}x_{j_{3}}.$
3. Multiplicative polynomials
Let $D$ be a positive integer and let $1\leq D<q\leq\infty$. Let
$P(\mathbf{x})=\sum_{k=1}^{D}\sum_{\Delta\in\mathcal{D}_{k}}\sum_{J\in\mathbf{N}^{k}}a_{\Delta,J}\ x_{J}^{\Delta}$
be a polynomial of degree at most $D$ in infinitely many variables $\mathbf{x}=(x_{j})_{j=1}^{\infty}$ with coefficients $a_{\Delta,J}$. We consider polynomials whose coefficients $a_{\Delta,J}$ are multiplicative in the following sense: For all $d\in\{1,2,\ldots,D\}$ there is a sequence
$\mathbf{a}_{d}=\left(a_{d,j}\right)_{j=1}^{\infty}\in\ell^{q/(q-d)}$
such that, if
$\Delta=(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}\operatorname{\qquad\text{and}\qquad}J=(j_{1},j_{2},\ldots,j_{k})\in\mathbf{N}^{k}$
then
$a_{\Delta,J}=a_{(d_{1},d_{2},\ldots,d_{k}),(j_{1},j_{2},\ldots,j_{k})}=a_{d_{1},j_{1}}a_{d_{2},j_{2}}\cdots a_{d_{k},j_{k}}.$
This gives the monomial factorization
$a_{\Delta,J}x_{J}^{\Delta}$ $=\left(a_{d_{1},j_{1}}a_{d_{2},j_{2}}\cdots a_{d_{k},j_{k}}\right)\left(x_{j_{1}}^{d_{1}}x_{j_{2}}^{d_{2}}\cdots x_{j_{k}}^{d_{k}}\right)$
$=\left(a_{d_{1},j_{1}}x_{j_{1}}^{d_{1}}\right)\left(a_{d_{2},j_{2}}x_{j_{2}}^{d_{2}}\right)\cdots\left(a_{d_{k},j_{k}}x_{j_{k}}^{d_{k}}\right)$
and so
$\sum_{J\in\mathbf{N}^{k}}a_{\Delta,J}\ x_{J}^{\Delta}$ $=\sum_{(j_{1},j_{2},\ldots,j_{j})\in\mathbf{N}^{k}}\left(a_{d_{1},j_{1}}x_{j_{1}}^{d_{1}}\right)\left(a_{d_{2},j_{2}}x_{j_{2}}^{d_{2}}\right)\cdots\left(a_{d_{k},j_{k}}x_{j_{k}}^{d_{k}}\right)$
$=\left(\sum_{j_{1}=1}^{\infty}a_{d_{1},j_{1}}x_{j_{1}}^{d_{1}}\right)\left(\sum_{j_{2}=1}^{\infty}a_{d_{2},j_{2}}x_{j_{2}}^{d_{2}}\right)\cdots\left(\sum_{j_{k}=1}^{\infty}a_{d_{k},j_{k}}x_{j_{k}}^{d_{k}}\right)$
$=\left(\mathbf{a}_{d_{1}},\mathbf{x}^{d_{1}}\right)\left(\mathbf{a}_{d_{2}},\mathbf{x}^{d_{2}}\right)\cdots\left(\mathbf{a}_{d_{k}},\mathbf{x}^{d_{k}}\right).$
For $\mathbf{x}=(x_{j})_{j=1}^{\infty}\in\ell^{q}$, the rearrangement is justified by the absolute convergence of the $k$ infinite series $\left(\mathbf{a}_{d_{1}},\mathbf{x}^{d_{1}}\right),\left(\mathbf{a}_{d_{2}},\mathbf{x}^{d_{2}}\right),\ldots,\left(\mathbf{a}_{d_{k}},\mathbf{x}^{d_{k}}\right)$. We obtain
$P(\mathbf{x})$ $=\sum_{k=1}^{D}\sum_{\Delta\in\mathcal{D}_{k}}\sum_{J\in\mathbf{N}^{k}}a_{\Delta,J}\ x_{J}^{\Delta}$
$=\sum_{k=1}^{D}\sum_{(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}}\left(\sum_{j_{1}=1}^{\infty}a_{d_{1},j_{1}}x_{j_{1}}^{d_{1}}\right)\left(\sum_{j_{2}=1}^{\infty}a_{d_{2},j_{2}}x_{j_{2}}^{d_{2}}\right)\cdots\left(\sum_{j_{k}=1}^{\infty}a_{d_{k},j_{k}}x_{j_{k}}^{d_{k}}\right)$
$=\sum_{k=1}^{D}\sum_{(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}}\left(\mathbf{a}_{d_{1}},\mathbf{x}^{d_{1}}\right)\left(\mathbf{a}_{d_{2}},\mathbf{x}^{d_{2}}\right)\cdots\left(\mathbf{a}_{d_{k}},\mathbf{x}^{d_{k}}\right).$
This is a finite sum, and so the series $P(\mathbf{x})$ converges absolutely for all $\mathbf{x}\in\ell^{q}$.
Let $D$ be a positive integer and let $1\leq D<q\leq\infty$. A $q$-multiplicative polynomial of degree at most $D$ or, simply, a multiplicative polynomial, is a polynomial of degree at most $D$ in infinitely many variables of the form
$P(\mathbf{x})=\sum_{k=1}^{D}\sum_{(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}}\left(\mathbf{a}_{d_{1}},\mathbf{x}^{d_{1}}\right)\left(\mathbf{a}_{d_{2}},\mathbf{x}^{d_{2}}\right)\cdots\left(\mathbf{a}_{d_{k}},\mathbf{x}^{d_{k}}\right)$
where
$\mathbf{a}_{d}=\left(a_{d,j}\right)_{j=1}^{\infty}\in\ell^{q/(q-d)}$
for all $d\in\{1,\ldots,D\}$. The infinite series converges for all $\mathbf{x}\in\ell^{q}$.
A multiplicative polynomial equation is an equation of the form
$P(\mathbf{x})=b$
where $P(\mathbf{x})$ is a multiplicative polynomial. We ask if this equation has a solution $\mathbf{x}\in\ell^{q}$.
The analogue of Theorem 1 for linear equations is the following “finitely many implies infinitely many” solvability result for multiplicative polynomial equations.
Theorem 3.
Let $D$ be a positive integer and let $1\leq D<q\leq\infty$. Let $I$ be an infinite set. For all $i\in I$ and $d\in\{1,2,\ldots,D\}$, let
$\mathbf{a}_{i,d}=\left(a_{i,d,j}\right)_{j=1}^{\infty}\in\ell^{q/(q-d)}.$
For $\Delta=(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}$ and $J=(j_{1},j_{2},\ldots,j_{k})\in\mathbf{N}^{k}$, let
$a_{i,\Delta,J}=a_{i,d_{1},j_{1}}a_{i,d_{2},j_{2}}\cdots a_{i,d_{k},j_{k}}.$
For all $i\in I$, consider the multiplicative polynomial equation
$P_{i}(\mathbf{x})$ $=\sum_{k=1}^{D}\sum_{\Delta\in\mathcal{D}_{k}}\sum_{J\in\mathbf{N}^{k}}a_{i,\Delta,J}\ x_{J}^{\Delta}$
$=\sum_{k=1}^{D}\sum_{(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}}\left(\mathbf{a}_{i,d_{1}},\mathbf{x}^{d_{1}}\right)\left(\mathbf{a}_{i,d_{2}},\mathbf{x}^{d_{2}}\right)\cdots\left(\mathbf{a}_{i,d_{k}},\mathbf{x}^{d_{k}}\right)$
$=b_{i}.$
Let $M>0$. If, for every finite subset $S$ of I, the finite set of polynomial equations $\{P_{i}(\mathbf{x})=b_{i}:i\in S\}$ has a solution $\mathbf{x}_{S}\in\ell^{q}$ with $\|\mathbf{x}_{S}\|_{q}\leq M$, then the infinite set of polynomial equations $\{P_{i}(\mathbf{x})=b_{i}:i\in I\}$ has a solution $\mathbf{x}\in\ell^{q}$ with $\|\mathbf{x}\|_{q}\leq M$.
4. Approximate finite implies exact infinite
The set $\{P_{i}(\mathbf{x})=b_{i}:i\in S\}$ of polynomial equations in infinitely many variables $\mathbf{x}=(x_{j})_{j=1}^{\infty}$ has an approximate solution if, for every $\varepsilon>0$, there exists a sequence $\mathbf{x}_{\varepsilon}$ such that $|P_{i}(\mathbf{x}_{\varepsilon})-b_{i}|\leq\varepsilon$ for all $i\in S$. Theorem 4 states that an infinite set of multiplicative polynomial equations has an exact solution if every finite subset of the equations has a norm-bounded approximate solution. This is the main result of this paper and immediately implies Theorems 1, 2, and 3.
The theorem and the lemmas in this section are valid in both the real and complex fields.
Theorem 4.
Let $D$ be a positive integer and let $1\leq D<q\leq\infty$. For all $k\in\{1,2,\ldots,D\}$, let
$\mathcal{D}_{k}=\left\{(d_{1},\ldots,d_{k})\in\mathbf{N}^{k}:d_{1}+d_{2}+\cdots+d_{k}\leq D\right\}.$
Let $I$ be an infinite set and let $\mathbf{x}=(x_{j})_{j=1}^{\infty}$. For all $i\in I$ and $d\in\{1,2,\ldots,D\}$, let
$\mathbf{a}_{i,d}=\left(a_{i,d,j}\right)_{j=1}^{\infty}\in\ell^{q/(q-d)}$
and
$\left(\mathbf{a}_{i,d},\mathbf{x}^{d}\right)=\sum_{j=1}^{\infty}a_{i,d,j}x_{j}^{d}.$
For $\Delta=(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}$ and $J=(j_{1},j_{2},\ldots,j_{k})\in\mathbf{N}^{k}$, define the multiplicative coefficient
$a_{i,\Delta,J}=a_{i,d_{1},j_{1}}a_{i,d_{2},j_{2}}\cdots a_{i,d_{k},j_{k}}.$
For all $i\in I$, the finite set of sequences $\{\mathbf{a}_{i,d}\}_{d=1}^{D}$ determines the multiplicative polynomial
$P_{i}(\mathbf{x})$ $=\sum_{k=1}^{D}\sum_{\Delta\in\mathcal{D}_{k}}\sum_{J\in\mathbf{N}^{k}}a_{i,\Delta,J}\ x_{J}^{\Delta}$
$=\sum_{k=1}^{D}\sum_{(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}}\left(\mathbf{a}_{i,d_{1}},\mathbf{x}^{d_{1}}\right)\left(\mathbf{a}_{i,d_{2}},\mathbf{x}^{d_{2}}\right)\cdots\left(\mathbf{a}_{i,d_{k}},\mathbf{x}^{d_{k}}\right).$
Let $M>0$ and $\mathbf{b}=(b_{i})_{i\in I}$. If, for every $\varepsilon>0$ and every finite subset $S$ of $I$, the finite set of polynomial inequalities
$\left\{|P_{i}(\mathbf{x})-b_{i}|\leq\varepsilon:i\in S\right\}$
has a solution $\mathbf{x}_{S,\varepsilon}\in\ell^{q}$ with $\|\mathbf{x}_{S,\varepsilon}\|_{q}\leq M$, then the infinite set of polynomial equations
$\left\{P_{i}(\mathbf{x})=b_{i}:i\in I\right\}$
has an exact solution $\mathbf{x}\in\ell^{q}$ with $\|\mathbf{x}\|_{q}\leq M$.
We begin with two results whose statements and proofs are valid in both the real and complex cases.
For $M>0$, the closed interval $[-M,M]=\{x\in\mathbf{R}:|x|\leq M\}$ and the closed ball $B_{M}=\{x\in\mathbf{C}:|x|\leq M\}$ are compact. For polynomial equations in $\mathbf{R}$ we use the compact topological space
$\Omega=\prod_{j=1}^{\infty}[-M,M]$
and for polynomial equations in $\mathbf{C}$ we use the compact topological space
$\Omega=\prod_{j=1}^{\infty}B_{M}.$
Let
$X_{q,M}=\left\{\mathbf{x}\in\ell^{q}:\|\mathbf{x}\|_{q}\leq M\right\}.$
Lemma 1.
The set $X_{q,M}$ is a compact subset of the topological space $\Omega$.
Proof.
If $\mathbf{x}=(x_{j})_{j=1}^{\infty}\in\ell^{q}$ and $\|\mathbf{x}\|_{q}\leq M$, then $|x_{j}|\leq M$ for all $j\in\mathbf{N}$ and so $X_{q,M}\subseteq\Omega$. Because $\Omega$ is compact, it suffices to prove that $X_{q,M}$ is closed in the product topology on $\Omega$, or equivalently, that the complement of $X_{q,M}$ is open. The complement of $X_{q,M}$ is the set
$Y_{q,M}$ $=\Omega\setminus X_{q,M}$
=$=\left\{\mathbf{y}\in\Omega:\mathbf{y}\notin\ell^{q}\right\}\bigcup\left\{\mathbf{y}\in\Omega:\mathbf{y}\in\ell^{q}\text{ and }\|\mathbf{y}\|_{q}>M\right\}$
=$=\left\{\mathbf{y}=(y_{j})_{j=1}^{\infty}:\sum_{j=1}^{\infty}|y_{j}|^{q}>M^{q}\right\}.$
Let $\mathbf{y}=(y_{j})_{j=1}^{\infty}\in Y_{q,M}$. There exist $\varepsilon>0$ and $N\in\mathbf{N}$ such that
$\sum_{j=1}^{N}|y_{j}|^{q}>M^{q}+\varepsilon.$
Because the function $f(t)=|t|^{q}$ is continuous, there exists $\delta>0$ such that $|t-y_{j}|<\delta$ implies
$|t|^{q}>|y_{j}|^{q}-\frac{\varepsilon}{2N}$
for all $j\in\{1,2,\ldots,N\}$. The set
$U=\left\{\mathbf{z}=(z_{j})_{j=1}^{\infty}\in\Omega:|z_{j}-y_{j}|<\delta\text{for all }j\in\{1,2,\ldots,N\}\right\}$
is an open neighborhood of $\mathbf{y}$ in $\Omega$. For all $\mathbf{z}=(z_{j})_{j=1}^{\infty}\in U$ we have
$\sum_{j=1}^{\infty}|z_{j}|^{q}$ $\geq\sum_{j=1}^{N}|z_{j}|^{q}\ >\sum_{j=1}^{N}\left(|y_{j}|^{q}-\frac{\varepsilon}{2N}\right)$
$=\sum_{j=1}^{N}|y_{j}|^{q}-\frac{\varepsilon}{2}>M^{q}+\frac{\varepsilon}{2}$
$>M^{q}$
and so $\mathbf{z}\in Y_{q,M}$. It follows that $U\subseteq Y_{q,M}$ and so $Y_{q,M}$ is an open subset of $\Omega$. This completes the proof. ∎
Lemma 2.
Let $d$ be a positive integer, $d<q\leq\infty$, and let $p=q/(q-d)$. Let $\mathbf{a}=(a_{j})_{j=1}^{\infty}\in\ell^{p}$. Let $M>0$. The function $f_{d}$ on $X_{q,M}$ defined by
$f_{d}(\mathbf{x})=(\mathbf{a},\mathbf{x}^{d})=\sum_{j=1}^{\infty}a_{j}x_{j}^{d}$
is continuous with respect to the product topology on $X_{q,M}$ as a subspace of $\Omega$.
Proof.
If $\mathbf{x}\in X_{q,M}$, then $\mathbf{x}^{d}\in\ell^{q/d}$. Because $(p,q/d)$ is a conjugate pair, the infinite series $\sum_{j=1}^{\infty}a_{j}x_{j}^{d}$ converges absolutely and the function $f_{d}(\mathbf{x})$ is well-defined.
Let $U$ be an open subset of $\mathbf{R}$ or $\mathbf{C}$. We shall prove $f_{d}^{-1}(U)$ is open in $X_{q,M}$ as a subspace of $\Omega$.
Let $\mathbf{x}=(x_{j})_{j=1}^{\infty}\in f_{d}^{-1}(U)$. Because $f_{d}(\mathbf{x})\in U$, there exists $\varepsilon>0$ such that $U$ contains the open set
$\left\{t:|t-f_{d}(\mathbf{x})|<\varepsilon\right\}.$
Because $\mathbf{a}\in\ell^{p}$, the series $\sum_{j=1}^{\infty}|a_{j}|^{p}$ converges and there is an integer $N_{\mathbf{a}}$ such that
$\sum_{j=N+1}^{\infty}|a_{j}|^{p}<\left(\frac{\varepsilon}{3M}\right)^{p}$
for all $N\geq N_{\mathbf{a}}$. Because $\mathbf{x}\in\ell^{q}$, the series $\sum_{j=1}^{\infty}\left|x_{j}\right|^{q}$ converges and there is an integer $N_{\mathbf{x}}$ such that
$\sum_{j=N+1}^{\infty}|x_{j}|^{q}<\left(\frac{\varepsilon}{3M}\right)^{q}$
for all $N\geq N_{\mathbf{x}}$. Choose $N\geq\max(N_{\mathbf{a}},N_{\mathbf{x}})$ and let $\delta>0$ satisfy
$\delta\sum_{j=1}^{N}|a_{j}|<\frac{\varepsilon}{3}.$
Choose $\delta^{\prime}>0$ such that $|y_{j}-x_{j}|<\delta^{\prime}$ implies $|y_{j}^{d}-x_{j}^{d}|<\delta$ for all $j\in\{1,2,\ldots,N\}$. The set
$V^{\prime}=\left\{\mathbf{y}=(y_{j})_{j=1}^{\infty}\in\Omega:|y_{j}-x_{j}|<\delta^{\prime}\text{ for all }j\in\{1,2,\ldots,N\}\right\}$
is an open neighborhood of $\mathbf{x}$ in the topological space $\Omega$ and so
$V$ $=V^{\prime}\cap X_{q,M}$
=$=\left\{\mathbf{y}=(y_{j})_{j=1}^{\infty}\in X_{q,M}:|y_{j}-x_{j}|<\delta^{\prime}\text{ for all }j\in\{1,2,\ldots,N\}\right\}$
is an open neighborhood of $\mathbf{x}$ in $X_{q,M}$. We shall prove that $V\subseteq f_{d}^{-1}(U)$.
Let $\mathbf{y}=(y_{j})_{j=1}^{\infty}\in V$. We define
$\tilde{x_{j}}=\begin{cases}0&\text{if $j\leq N$}\\ x_{j}&\text{if $j\geq N+1$}\end{cases}$
and
$\tilde{y_{j}}=\begin{cases}0&\text{if $j\leq N$}\\ y_{j}&\text{if $j\geq N+1$.}\end{cases}$
Then
$\tilde{\mathbf{x}}=(\tilde{x_{j}})_{j=1}^{\infty}\in\ell^{q}\operatorname{\qquad\text{and}\qquad}\tilde{\mathbf{y}}=(\tilde{y_{j}})_{j=1}^{\infty}\in\ell^{q}$
and
$\|\tilde{\mathbf{x}}\|_{q}\leq\|\mathbf{x}\|_{q}\leq M\operatorname{\qquad\text{and}\qquad}\|\tilde{\mathbf{y}}\|_{q}\leq\|\mathbf{y}\|_{q}\leq M.$
Thus, $\tilde{\mathbf{x}}\in X_{q,M}$ and $\tilde{\mathbf{y}}\in X_{q,M}$. We define
$\tilde{a_{j}}=\begin{cases}0&\text{if $j\leq N$}\\ a_{j}&\text{if $j\geq N+1$.}\end{cases}$
Then
$\tilde{\mathbf{a}}=(\tilde{a_{j}})_{j=1}^{\infty}\in\ell^{p}$
and
$\|\tilde{\mathbf{a}}\|_{p}\leq\frac{\varepsilon}{3M}.$
We have
$f_{d}(\mathbf{y})-f_{d}(\mathbf{x})$ $=(\mathbf{a},\mathbf{y}^{d})-(\mathbf{a},\mathbf{x}^{d})$
$=\sum_{j=1}^{\infty}a_{j}y_{j}^{d}-\sum_{j=1}^{\infty}a_{j}x_{j}^{d}=\sum_{j=1}^{\infty}a_{j}\left(y_{j}^{d}-x_{j}^{d}\right)$
$=\sum_{j=1}^{N}a_{j}(y_{j}^{d}-x_{j}^{d})+\sum_{j=N+1}^{\infty}a_{j}(y_{j}^{d}-x_{j}^{d}).$
Applying the triangle, Hölder, and Minkowski inequalities, we obtain
$\left|f_{d}(\mathbf{y})-f_{d}(\mathbf{x})\right|$ $\leq\sum_{j=1}^{N}|a_{j}|\ |y_{j}^{d}-x_{j}^{d}|+\sum_{j=N+1}^{\infty}|a_{j}(y_{j}^{d}-x_{j}^{d})|$
$\leq\delta\sum_{j=1}^{N}|a_{j}|+\left\|\left(\tilde{\mathbf{a}},\tilde{\mathbf{y}}-\tilde{\mathbf{x}}\right)\right\|_{1}$
$<\frac{\varepsilon}{3}+\left\|\tilde{\mathbf{a}}\right\|_{p}\left\|\tilde{\mathbf{y}}-\tilde{\mathbf{x}}\right\|_{q}$
$\leq\frac{\varepsilon}{3}+\left\|\tilde{\mathbf{a}}\right\|_{p}\left(\left\|\tilde{\mathbf{y}}\right\|_{q}+\left\|\tilde{\mathbf{x}})\right\|_{q}\right)$
$<\frac{\varepsilon}{3}+\frac{\varepsilon}{3M}(M+M)$
$=\varepsilon.$
Therefore, $f_{d}(\mathbf{y})\in U$ and $V\subseteq f_{d}^{-1}(U)$. Thus, the set $f_{d}^{-1}(U)$ is open in $X_{q,M}$ and the function $f_{d}(\mathbf{x})=(\mathbf{a},\mathbf{x}^{d})$ is continuous on $X_{q,M}$. This completes the proof. ∎
We can now prove Theorem 4.
Proof.
For all $i\in I$ we have the multiplicative polynomial
$P_{i}(\mathbf{x})=\sum_{k=1}^{D}\sum_{(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}}\left(\mathbf{a}_{i,d_{1}},\mathbf{x}^{d_{1}}\right)\left(\mathbf{a}_{i,d_{2}},\mathbf{x}^{d_{2}}\right)\cdots\left(\mathbf{a}_{i,d_{k}},\mathbf{x}^{d_{k}}\right).$
By Lemma 1, the set $X_{q,M}$ is a compact subset of $\Omega$. By Lemma 2, the functions $\left(\mathbf{a}_{i,d_{i}},\mathbf{x}^{d_{i}}\right)$ are continuous on $X_{q,M}$ for all $d_{i}\in\{1,\ldots,D\}$. Finite sums of finite products of continuous functions are continuous, and so the multiplicative polynomials $P_{i}(\mathbf{x})$ are continuous functions on $X_{q,M}$ for all $i\in I$. It follows that, for all $i\in I$ and $\varepsilon>0$, the approximation set
$F_{i,\varepsilon}=\left\{\mathbf{x}\in X_{q,M}:|P_{i}(\mathbf{x})-b_{i}|\leq\varepsilon\right\}$
is a closed subset of the compact set $X_{q,M}$. For every finite subset $S$ of $I$, the set of polynomial inequalities
$\left\{|P_{i}(\mathbf{x})-b_{i}|\leq\varepsilon:i\in S\right\}$
has a solution $\mathbf{x}_{S,\varepsilon}\in\ell^{q}$ with $\|\mathbf{x}_{S,\varepsilon}\|_{q}\leq M$ and so the set of closed sets $\{F_{i,\varepsilon}:i\in I\text{ and }\varepsilon>0\}$ has the finite intersection property. Therefore,
$\bigcap_{\begin{subarray}{c}i\in I\\ \varepsilon>0\end{subarray}}F_{i,\varepsilon}=\left\{\mathbf{x}\in X_{q,M}:P_{i}(\mathbf{x})=b_{i}\text{ for all $i\in I$}\right\}$
is nonempty and the infinite set of polynomial equations has an exact solution. This completes the proof Theorem 4. ∎
Theorem 2 for infinitely many linear equations in infinitely many variables is the special case of Theorem 4 with $D=1$. Theorem 4 implies Theorem 3 and Theorem 2 implies Theorem 1.
We have the following refinement of Theorem 4.
Theorem 5.
With the hypotheses of Theorem 4, let
$\mathbf{m}=(m_{j})_{j=1}^{\infty}\in\ell^{q}\qquad\text{ with $m_{j}\geq 0$ for all $j\in\mathbf{N}$. }$
If, for every $\varepsilon>0$ and every finite subset $S$ of I, the finite set of polynomial inequalities
$\left\{|P_{i}(\mathbf{x})-b_{i}|\leq\varepsilon:i\in S\right\}$
has a solution $\mathbf{x}_{S,\varepsilon}=\left(x_{S,\varepsilon,j}\right)_{j=1}^{\infty}\in\ell^{q}$ with $|x_{S,\varepsilon,j}|\leq m_{j}$ for all $j\in\mathbf{N}$, then the infinite set of polynomial equations
$\left\{P_{i}(\mathbf{x})=b_{i}:i\in I\right\}$
has an exact solution $\mathbf{x}=\left(x_{j}\right)_{j=1}^{\infty}\in\ell^{q}$ with $|x_{j}|\leq m_{j}$ for all $j\in\mathbf{N}$.
Proof.
Let $M=\|\mathbf{m}\|_{q}$. For polynomial equations in $\mathbf{R}$ we use the compact space
$X_{\mathbf{m}}=\prod_{j=1}^{\infty}[-m_{j},m_{j}]$
and for polynomial equations in $\mathbf{C}$ we use the compact space
$X_{\mathbf{m}}=\prod_{j=1}^{\infty}B_{m_{j}}.$
If $\mathbf{x}\in X_{\mathbf{m}}$, then $\mathbf{x}\in\ell^{q}$ and $\|\mathbf{x}\|_{q}\leq M$, and so
$X_{\mathbf{m}}\subseteq X_{q,M}=\left\{\mathbf{x}\in\ell^{q}:\|\mathbf{x}\|_{q}\leq M\right\}.$
By Lemma 1, the set $X_{q,M}$ is a compact subset of $\Omega$. By Lemma 2, for all $d\in\{1,\ldots,D\}$ and for all $\mathbf{a}\in\ell^{q/(q-d)}$, the function $f_{d}(\mathbf{x})=(\mathbf{a},\mathbf{x}^{d})$ is continuous on $X_{q,M}$ and so its restriction to the compact set $X_{\mathbf{m}}$ is continuous. It follows that the multiplicative polynomials $P_{i}(\mathbf{x})$ are continuous functions on $X_{\mathbf{m}}$ for all $i\in I$, and so the approximation set
$F_{i,\varepsilon}=\left\{\mathbf{x}\in X_{\mathbf{m}}:|P_{i}(\mathbf{x})-b_{i}|\leq\varepsilon\right\}$
is a nonempty closed subset of the compact set $X_{\mathbf{m}}$. These sets have the finite intersection property and so
$\bigcap_{\begin{subarray}{c}i\in I\\ \varepsilon>0\end{subarray}}F_{i,\varepsilon}=\left\{\mathbf{x}\in X_{\mathbf{m}}:P_{i}(\mathbf{x})=b_{i}\text{ for all $i\in I$}\right\}$
is nonempty. This completes the proof ∎
5. An application to Dirichlet series
Here is an example of a multiplicative polynomial. Let $D$ be a positive integer and let $D<q\leq\infty$. For $d\in\{1,\ldots,D\}$, let ${s_{d}}\in\mathbf{C}$ with $\Re({s_{d}})>(q-d)/q$. Then
$\Re\left(\frac{qs_{d}}{q-d}\right)>1$
and so the Riemann zeta function
$\zeta\left(\frac{q{s_{d}}}{q-d}\right)=\sum_{j=1}^{\infty}\frac{1}{j^{q{s_{d}}/(q-d)}}=\sum_{j=1}^{\infty}\left(\frac{1}{j^{s_{d}}}\right)^{q/(q-d)}$
converges absolutely. Consider the sequence
$\mathbf{a}_{d}=\left(a_{d,j}\right)_{j=1}^{\infty}=\left(\frac{1}{j^{s_{d}}}\right)_{j=1}^{\infty}\in\ell^{q/(q-d)}$
for all $d\in\{1,\ldots,D\}$, where
$a_{d,j}=\frac{1}{j^{s_{d}}}$
and the polynomial
$(\mathbf{a}_{d},\mathbf{x}^{d})=\sum_{j=1}^{\infty}a_{d,j}x_{j}^{d}=\sum_{j=1}^{\infty}\frac{x_{j}^{d}}{j^{s_{d}}}.$
Let $k\in\{1,\ldots,D\}$. For all
$\Delta=(d_{1},d_{2},\ldots,d_{k})\in\mathcal{D}_{k}\operatorname{\qquad\text{and}\qquad}J=(j_{1},j_{2},\ldots,j_{k})\in\mathbf{N}^{k}$
we define the multiplicative coefficient
$a_{\Delta,J}$ $=a_{(d_{1},d_{2},\ldots,d_{k}),(j_{1},j_{2},\ldots,j_{k})}=a_{d_{1},j_{1}}a_{d_{2},j_{2}}\cdots a_{d_{k},j_{k}}$
$=\frac{1}{j_{1}^{s_{d_{1}}}\cdots j_{k}^{s_{d_{k}}}}$
. Then
$\sum_{J\in\mathbf{N}^{k}}a_{\Delta,J}\ x_{J}^{\Delta}$ $=\left(\mathbf{a}_{d_{1}},\mathbf{x}^{d_{1}}\right)\left(\mathbf{a}_{d_{2}},\mathbf{x}^{d_{2}}\right)\cdots\left(\mathbf{a}_{d_{k}},\mathbf{x}^{d_{k}}\right)$
$=\sum_{(j_{1},\ldots,j_{k})\in\mathbf{N}^{k}}\frac{x_{j_{1}}^{d_{1}}\cdots x_{j_{k}}^{d_{k}}}{j_{1}^{s_{1}}\cdots j_{k}^{s_{k}}}.$
We obtain the $q$-multiplicative polynomial of degree at most $D$
$P(\mathbf{x})$ $=P(\mathbf{x},s_{1},\ldots,s_{D})$
$=\sum_{k=1}^{D}\sum_{\Delta\in\mathcal{D}_{k}}\sum_{J\in\mathbf{N}^{k}}a_{\Delta,J}\ x_{J}^{\Delta}$
$=\sum_{k=1}^{D}\sum_{(d_{1},\ldots,d_{k})\in\mathcal{D}_{k}}\sum_{(j_{1},\ldots,j_{k})\in\mathbf{N}^{k}}\frac{x_{j_{1}}^{d_{1}}\cdots x_{j_{k}}^{d_{k}}}{j_{1}^{s_{1}}\cdots j_{k}^{s_{k}}}.$
This series converges absolutely for all $\mathbf{x}=(x_{j})_{j=1}^{\infty}\in\ell^{q}$.
For $D=1$ and $q=2$ and for complex numbers $s$ with $\Re(s)>1/2$, the multiplicative polynomial is the Dirichlet series
$L(\mathbf{x})=\sum_{n=1}^{\infty}\frac{x_{n}}{n^{s}}.$
This series converges absolutely for all $\mathbf{x}=(x_{n})_{n=1}^{\infty}\in\ell^{2}$. The corresponding linear equation in infinitely many variables is simply
$\sum_{n=1}^{\infty}\frac{x_{n}}{n^{s}}=b.$
Applying Theorem 2, we obtain the following result, which asserts the existence of a Dirichlet series taking prescribed values at a given infinite set of points in the half-plane $\Re(s)>1/2$.
Theorem 6.
Let $I$ be an infinite set and let $(b_{i})_{i\in I}$ and $(s_{i})_{i\in I}$ be sequences of complex numbers with $\Re(s_{i})>1/2$ for all $i\in I$. Consider the linear equation in infinitely many variables
$L_{i}(\mathbf{x})=\sum_{n=1}^{\infty}\frac{x_{n}}{n^{s_{i}}}=b_{i}.$
Let $M>0$. If, for every finite subset $S$ of I and every $\varepsilon>0$, the finite set of linear inequalities
$\left|\sum_{n=1}^{\infty}\frac{x_{n}}{n^{s_{i}}}-b_{i}\right|\leq\varepsilon\qquad\text{for all $i\in S$}$
has a solution $\mathbf{x}_{S,\varepsilon}\in\ell^{2}$ with $\|\mathbf{x}_{S,\varepsilon}\|_{2}\leq M$, then the infinite set of linear equations
$\sum_{n=1}^{\infty}\frac{x_{n}}{n^{s_{i}}}=b_{i}\qquad\text{for all $i\in I$}$
has an exact solution $\mathbf{x}\in\ell^{2}$ with $\|\mathbf{x}\|_{2}\leq M$.
6. Open problems
. The results in this paper suggest several questions.
  1. (1)
    Are Theorems 34, and 5 true for infinite sets of polynomial equations in which the polynomials are not multiplicative?
  2. (2)
    Are there classes $\mathcal{C}$ of infinite sets of polynomial equations in infinitely many variables for which there is an integer $S=S(\mathcal{C})$ such that, if every finite set of at most $S$ equations has a solution or an approximate solution, then the infinite set of equations has a solution?
  3. (3)
    Let $E$ be a subfield of the complex numbers. Let $\{P_{i}(\mathbf{x})=b_{i}:i\in I\}$ be an infinite set of polynomial equations in which every finite subset of the equations has an exact solution or an approximate solution in $E$.
    1. (a)
      For what subfields $E$ of the complex numbers is it true that the infinite set of equations have an exact solution in $E$?
    2. (b)
      For what sets of polynomial equations might this be true?
  4. (4)
    The “finitely many implies infinitely many” paradigm applies to linear equations and multiplicative polynomial equations with bounded or sequentially bounded norms (Theorems 34, and 5). For equations of this special type, countably many implies finitely many. But there may be other classes of equations for which this paradigm does not hold, but a stronger condition (“countably many implies uncountably many”) is true. The problem is to determine if such “new” classes of equations exist.
  5. (5)
    Is there an uncountably infinite set of equations such that every countably infinite subset of the equations has an exact or an approximate solution, but the uncountably infinite set of equations has no exact solution?
Appendix A The theorems of F. Riesz and Abian-Eslami
Theorem 7.
Let $(p,q)$ be a conjugate pair and let $M>0$. Let $\mathbf{a}_{i}=(a_{i,j})_{j=1}^{\infty}\in\ell^{p}$ for all $i\in\mathbf{N}$ and let $(b_{i})_{i\in\mathbf{N}}$ be a sequence of numbers. The following are equivalent.
  1. (a)
    For all $r\in\mathbf{N}$, there exists $\mathbf{x}_{r}=(x_{r,j})_{j=1}^{\infty}\in\ell^{q}$ such that
    $\|\mathbf{x}_{r}\|_{q}\leq M\operatorname{\qquad\text{and}\qquad}\sum_{j=1}^{\infty}a_{i,j}x_{r,j}=b_{i}\qquad\text{for all $i\in\{1,2,3,\ldots,r\}$.}$
  2. (b)
    For all $r\in\mathbf{N}$ and $h_{1},\ldots,h_{r}\in\mathbf{R}$,
    $\left|\sum_{i=1}^{r}h_{i}b_{i}\right|\leq M\left(\sum_{j=1}^{\infty}\left|\sum_{i=1}^{r}h_{i}a_{i,j}\right|^{p}\right)^{1/p}.$
  3. (c)
    There exists $\mathbf{x}=(x_{j})_{j=1}^{\infty}\in\ell^{q}$ such that
    $\|\mathbf{x}\|_{q}\leq M\operatorname{\qquad\text{and}\qquad}\sum_{j=1}^{\infty}a_{i,j}x_{j}=b_{i}\qquad\text{for all $i\in\mathbf{N}$.}$
F. Riesz [5] proved the equivalence of (b) and (c) and Abian-Eslami [1] proved (a) implies (c).
Proof.
First we prove that (a) implies (b). For all $h_{1},\ldots,h_{r}$,
$\sum_{i=1}^{r}h_{i}\mathbf{a}_{i}=\sum_{i=1}^{r}h_{i}(a_{i,j})_{j=1}^{\infty}=\left(\sum_{i=1}^{r}h_{i}a_{i,j}\right)_{j=1}^{\infty}\in\ell^{p}$
and
$\left\|\sum_{i=1}^{r}h_{i}\mathbf{a}_{i}\right\|_{p}=\left(\sum_{j=1}^{\infty}\left|\sum_{i=1}^{r}h_{i}a_{i,j}\right|^{p}\right)^{1/p}.$
We have $\mathbf{x}_{r}=(x_{r,j})_{j=1}^{\infty}\in\ell^{q}$ and so, by Hölder’s inequality,
$\left(\sum_{i=1}^{r}h_{i}\mathbf{a}_{i}\right)\mathbf{x}_{r}=\left(\sum_{i=1}^{r}h_{i}a_{i,j}x_{j}\right)_{j=1}^{\infty}\in\ell^{1}$
and
$\left\|\left(\sum_{i=1}^{r}h_{i}\mathbf{a}_{i}\right)\mathbf{x}_{r}\right\|_{1}=\sum_{j=1}^{\infty}\left|\sum_{i=1}^{r}h_{i}a_{i,j}x_{j}\right|<\infty.$
For every positive integer $N$, let
$g(N)=\sum_{j=N+1}^{\infty}\left|\sum_{i=1}^{r}h_{i}a_{i,j}x_{j}\right|.$
We have
$\lim_{N\rightarrow\infty}g(N)=0$
and
$\left|\sum_{i=1}^{r}h_{i}b_{i}\right|$ $=\left|\sum_{i=1}^{r}h_{i}\left(\sum_{j=1}^{\infty}a_{i,j}x_{r,j}\right)\right|$
$\leq\left|\sum_{i=1}^{r}h_{i}\sum_{j=1}^{N}a_{i,j}x_{r,j}\right|+\left|\sum_{i=1}^{r}h_{i}\sum_{j=N+1}^{\infty}a_{i,j}x_{r,j}\right|$
$\leq\left|\sum_{j=1}^{N}\sum_{i=1}^{r}h_{i}a_{i,j}x_{r,j}\right|+\sum_{j=N+1}^{\infty}\left|\sum_{i=1}^{r}h_{i}a_{i,j}x_{r,j}\right|$
$\leq\sum_{j=1}^{N}\left|\sum_{i=1}^{r}h_{i}a_{i,j}x_{r,j}\right|+g(N)$
$\leq\sum_{j=1}^{\infty}\left|\sum_{i=1}^{r}h_{i}a_{i,j}x_{r,j}\right|+g(N)$
$=\left\|\left(\sum_{i=1}^{r}h_{i}\mathbf{a}_{i}\right)\mathbf{x}_{r}\right\|_{1}+g(N)$
$\leq\left\|\sum_{i=1}^{r}h_{i}\mathbf{a}_{i}\right\|_{p}\|\mathbf{x}_{r}\|_{q}+g(N)$
$\leq M\left(\sum_{j=1}^{\infty}\left|\sum_{i=1}^{r}h_{i}a_{i,j}\right|^{p}\right)^{1/p}+g(N).$
This inequality is valid for all $N$ and $\lim_{N\rightarrow\infty}g(N)=0$. Therefore,
$\left|\sum_{i=1}^{r}h_{i}b_{i}\right|\leq M\left(\sum_{j=1}^{\infty}\left|\sum_{i=1}^{r}h_{i}a_{i,j}\right|^{p}\right)^{1/p}.$
Thus, (a) implies (b).
Next we prove that (b) implies (c). Let $W$ be the vector subspace of $\ell^{p}$ spanned by the set $\{\mathbf{a}_{i}:i\in\mathbf{N}\}$. Let $h_{1},\ldots,h_{r},h^{\prime}_{1},\ldots,h^{\prime}_{r}$ be scalars such that
$\sum_{i=1}^{r}h_{i}\mathbf{a}_{i}=\sum_{i=1}^{r}h^{\prime}_{i}\mathbf{a}_{i}.$
Then
$\sum_{i=1}^{r}(h_{i}-h^{\prime}_{i})\mathbf{a}_{i}=\mathbf{0}$
and
$\left|\sum_{i=1}^{r}h_{i}b_{i}-\sum_{i=1}^{r}h^{\prime}_{i}b_{i}\right|$ $=\left|\sum_{i=1}^{r}(h_{i}-h^{\prime}_{i})b_{i}\right|$
$\leq M\left(\sum_{j=1}^{\infty}\left|\sum_{i=1}^{r}(h_{i}-h^{\prime}_{i})a_{i,j}\right|^{p}\right)^{1/p}$
$=M\left\|\sum_{i=1}^{r}(h_{i}-h^{\prime}_{i})\mathbf{a}_{i}\right\|_{p}=0.$
It follows that there is a well-defined linear functional $f$ on $W$ such that
$f\left(\sum_{i=1}^{r}h_{i}\mathbf{a}_{i}\right)=\sum_{i=1}^{r}h_{i}b_{i}.$
In particular, $f_{d}(\mathbf{a}_{i})=b_{i}$. Because
$\left|f\left(\sum_{i=1}^{r}h_{i}\mathbf{a}_{i}\right)\right|=\left|\sum_{i=1}^{r}h_{i}b_{i}\right|\leq M\left\|\sum_{i=1}^{r}h_{i}\mathbf{a}_{i}\right\|_{p}$
for all $\sum_{i=1}^{r}h_{i}\mathbf{a}_{i}\in W$, the linear functional $f$ has norm $\|f\|\leq M$. By the Hahn-Banach theorem, there is a bounded linear functional $F$ on $\ell^{p}$ such that $F(\mathbf{w})=f_{d}(\mathbf{w})$ for all $\mathbf{w}\in W$ and $\|F\|\leq M$.
For every bounded linear functional $F$ on $\ell^{p}$ there is a sequence $\mathbf{x}\in\ell^{q}$ such that $F(\mathbf{a})=(\mathbf{a},\mathbf{x})$ for all $\mathbf{a}\in\ell^{p}$. For all $i\in\mathbf{N}$ we have
$b_{i}=f_{d}(\mathbf{a}_{i})=F(\mathbf{a}_{i})=(\mathbf{a}_{i},\mathbf{x})=\sum_{j=1}^{\infty}a_{i,j}x_{j}.$
Thus, (b) implies (c).
The proof that (c) implies (a) is immediate. ∎